OpenStudy (anonymous):

How old am I if 400 reduced by 2 times my age is 244? Give an equation

OpenStudy (s3a):

400-2*age=244 400-244 - 2*age = 244-244 156 - 2*age=0 156 - 2*age+2*age = 0+2*age 156 = 2*age 156/2 = 2*age/2 78 = age (You can replace age with the letter x.)